COMPLICATED THREEFOLD SYMMETRIC MOLECULAR SYSTEMS \underline{JOHN F. STANTON}, Institute for Theoretical Chemistry, Department of Chemistry, University of Texas at Austin, Austin, TX 78712.

Some results and insight into the interesting spectroscopy of the Si_{3} and NO_{3} molecules will be presented.